# Methodology, Software, Validation, Writing-Original Draft, Writing-Review and Editing, Visualization

### v1

@@ Abstract
Autism spectrum disorder (ASD) is a pervasive developmental disorder of the central nervous system, which occurs most frequently in childhood and is characterized by unusual and repetitive ritualistic behaviors. Currently, diagnostic methods primarily rely on questionnaire surveys and behavioral observation, which may lead to misdiagnoses due to the subjective evaluation and measurement used in these traditional methods. With the advancement in medical imaging, MR imaging-based diagnosis has become an alternative and more objective approach. In this paper, we propose a Hybrid neural Network model for ASD identification, termded ASD-HNet, to hierarchically extract features on the functional brain networks based on resting-state functional magnetic resonance imaging data. This hierarchical method can better extract brain representations, improve the diagnostic accuracy, and help us better locate brain regions related to ASD. Specifically, features are extracted from three scales: local regions of interest (ROIs) scale, community-clustering scale, and the whole-communities scale. For the ROI scale, graph convolution is used to transfer features between ROIs. At the community cluster scale, functional gradients are introduced, the clustering algorithm K-Means is used to automatically cluster ROIs with similar functional gradients into several communities, and features of ROIs belonging to the same community are extracted to characterize these communities. At global information integration scale, we extract global features from community-scale brain networks to characterize the whole brain networks. We validate the effectiveness of our method using the public dataset of Autism Brain Imaging Data Exchange I (ABIDE I), and elucidate the interpretability of the method. Experimental results demonstrate that the proposed ASD-HNet can yield superior performance than compared methods.
@@ 1 Introduction
Autism spectrum disorder (ASD) is a type of pervasive developmental disorder affecting the central nervous system (Pandolfi et al., 2018), approximately 1% of the world’s population is affected by ASD. Moreover, ASD is associated with a high incidence of concurrent disorders, such as depression and anxiety (Vohra et al., 2017). Autism generally occurs in childhood, and its impairment may remain lifelong, and parts of individuals with ASD require lifelong support (Lord et al., 2018). ASD encompasses a range of symptoms, such as social communication disorders and restrictive repetitive behaviors (Hirota and King, 2023; Bhat et al., 2014). Autism is extremely damaging to the physical and mental health of patients, and significantly affects the lives of affected individuals and their families. However, due to the lack of specific pathogenesis, it is very challenging to make an accurate diagnosis.
Unfortunately, there are no definitive drugs and methods for the treatment of ASD, and the pathological mechanisms of ASD remain for further exploring (Geschwind, 2009). Currently, the most effective approach is to provide advanced behavioral training for individuals with ASD, implement early intervention for ASD, and correct the problem behaviors of autistic patients. This necessitates the early and accurate diagnosis of individuals with ASD, which is crucial. However, by the time the individual exhibits the corresponding behaviors, the patient may miss the optimal treatment time, which makes treatment more difficult. At present, most diagnostic methods are still through questionnaires and behavioral observation to determine whether an individual has ASD (American Psychiatric Association et al., 2013). This method has strong subjectivity and is influenced by the environment, which can lead to a higher risk of misdiagnosis (Wang et al., 2022). There is an urgent need for a highly accurate pathological diagnostic method and the identification of more autism-related biomarkers to aid in diagnosis.
@@
Benefiting from the rapid development of medical image technology, imaging-based approaches provides us with important tools to analyze and study neurophysiology. Functional magnetic resonance imaging (fMRI), as a non-invasive neuroimaging technique (Matthews and Jezzard, 2004), can measure the blood oxygen level dependent (BOLD) changes caused by neuronal activity throughout the brain at a series of time points (Friston et al., 1994). Many studies have demonstrated the feasibility of using resting state functional magnetic resonance imaging (rs-fMRI) to reveal interactions between regions of interest (ROIs) in the brain of people with mental illness (Dvornek et al., 2018). In recent years, numerous researchers have integrated deep learning technology with rs-fMRI to conduct corresponding exploration and research on mental diseases, yielding significant achievements (Dong et al., 2020; Luppi et al., 2022). Most studies divide the brain into multiple ROIs using prefabricated brain templates, obtain the BOLD time series of ROIs, and then calculate the BOLD time series correlation between ROIs to obtain the brain functional connectivity network (FCN) for corresponding analysis and exploration of the brain (Dong and Sun, 2024; Wang et al., 2023; Hong et al., 2019).
@@
Previous deep learning models often lack the exploration and application of physiological mechanisms, as well as multi-scale features mining, e.g. from local regions to global brain scale, resulting in suboptimal diagnostic performance of the models. In current study, we combine rs-fMRI with deep learning techniques to design a network model from a brain physiology perspective, aiming to boost the detection performance on ASD. We propose a hybrid neural network model for ASD identification, termded ASD-HNet, to hierarchically extract features on the functional brain networks at three scales, i.e., local ROI scale, community-clustering scale, and whole-communities scale. Specifically, we use the AAL template (Tzourio-Mazoyer et al., 2002) to divide the brain into 116 ROIs, extract the BOLD signals in each ROI, and then calculate the functional connectivity network. Then, the graph convolutional network (GCN) (Niepert et al., 2016) is used to transfer information between ROIs. After that, we use functional gradients to automatically cluster ROIs into several communities instead of using another scale of prefabricated atlas. In this way, ROIs with similar functions are clustered into a community, allowing us to obtain the community-scale features. Finally, we perform global brain convolution on the community-scale features to integrate features on global scale. Furthermore, based on the features obtained by the ASD-HNet, we introduced prototype learning for classification to achieve better identification results. The experiments on the public dataset of Autism Brain Imaging Data Exchange I (ABIDE I) verified the superior performance of the proposed ASD-HNet.
@@ 2.1 Functional Gradients
The functional gradients of the brain can provide a more comprehensive perspective by capturing the continuous spatial patterns of connectivity outside the segregated networks in the human brain (Guell et al., 2018; Huntenburg et al., 2018). A gradient-based approach, which is a nonlinear decomposition of high-dimensional rest-state functional connectivity (rs-FC), can identify brain functional levels by representing brain connectivity in a continuous, low-dimensional space (Dong et al., 2020). Functional gradients have been widely applied in brain analysis (Dong et al., 2020; Guo et al., 2023; Gong and Zuo, 2023). In the analysis of various mental diseases, functional gradients can obviously reflect the abnormal functional hierarchy of brain regions.
@@ 2.2 Materials and data preprocessing
For data preprocessing, we used the data Processing Assistant for Resting-State fMRI (DPARSF) to preprocess rs-fMRI data (Yan and Zang, 2010). The specific steps are as follows: (1) Discarding the first 5 volumes. (2) Correcting of slice-related delay. (3) Correcting head movement. (4) Using EPI template normalization in MNI space, resampling to 3 mm resolution of $3\times 3\times 3$. (5) Smoothing with a 4 mm full-width at half-maximum (FWHM) Gaussian kernel. (6) Linear detrending and temporal band-pass filtering of BOLD signal (0.01-0.10 Hz). (7) Harmful signals that return head movement parameters, white matter, cerebrospinal fluid (CSF), and global signals. The registered fMRI volumes are then partitioned into 116 ROIs using the AAL template, and BOLD time series are obtained for each ROI. Based on these time series of all the ROIs, we obtained the brain FCN matrix by calculating the Pearson correlation coefficient with the formula (1).
@@ 2.3 Proposed Method
The schematic diagram of the proposed ASD-HNet is shown in Figure 2. This model is designed to progressively extract the discriminative and robust representations based on FCNs from three different scales, i.e., local ROI scale, community-clustering scale, and whole-communities scale, to achieve ASD identification. More details are described below.
@@ 2.3.1 The local ROI scale
With the $F$-score, we can obtain two sparse FCN matrices. We kept 10% of the connections in the functional connectivity matrix to create a sparse FCN matrix, which was used to create a graph network. Besides, we removing half of connections in the functional connectivity matrix to create another sparse FCN matrix, which was used as the node features in the graph network.
Based on these two sparse FCN matrices, we constructed a ROI-scale graph network $G(V,E)$, where $V$ is the node eigenmatrix and $E$ is the adjacency matrix. Then, we used graph convolutional neural network to conduct feature extraction on the ROI-level scale. Here, we used one layer of spatial-based GCN to complete the message transfer between ROIs, the specific implementation is as follows:
@@
where $W$ is a layer-specific trainable weight matrix with the shape (116, 64) and $\sigma$ is the activation function. In order to make nodes pay attention to their own characteristics, the diagonal elements of the adjacency matrix $E$ are all set to 1.
After graph convolution, we further employed a ROI-level attention mechanism to weight each ROI features. Sepcifically, we initialized an attention weight $H_{atten}$ with all values set to 1, and then used two successive fully connected layers to learn the attention weight for each ROI. The operation of one fully connected layer is shown in the following formula (4). It should be noted that since ROIs are weighted using positive numbers, so the activation function ($\sigma_{atten}$) was ReLU in the fully connected layer.
@@
After obtaining the attention weights of ROIs, we used these weights to multiple each ROI, and then used a residual connection to fuse the weighted and original ROI-level features, following by a batch normalization operation, as shown in formula (5).
@@
2.3.2 The community-clustering scale
The brain’s community network comprises groups of functionally similar ROIs, each with a high degree of functional concentration. Community networks are crucial for understanding the functional organization of the brain (Van Den Heuvel et al., 2009). Functional communities have been shown to be relevant to understanding cognitive behavior (Van Den Heuvel and Pol, 2010), mental states (Geerligs et al., 2015), and neurological and psychiatric disorders (Canario et al., 2021). We used second module further extracted functional features at community-clustering scale as shown in Figure 2. Since functional gradients can well reflect the functional similarity between ROIs, we proposed to use functional gradients to automatically cluster ROIs into community networks.
We utilized whole group-level (include ASD subjects and HC subjects) average functional connectivity matrix inputs into BrainSpace (Vos de Wael et al., 2020) to obtain group-level functional gradients. In general, the calculation of functional gradients is mainly divided into two steps. The first step is to sparse the input matrix (the sparsity is 10%), and then select a kernel function to calculate the affinity matrix. In the second step, the functional gradients are obtained by using nonlinear dimensionality reduction on the resulting affinity matrix. In our work, we chose cosine similarity as kernel function to calculate the affinity matrix ($A(i,j)$). The calculation formula is as follows:
@@ 2.3.2 The community-clustering scale
where $A$ is the affinity matrix, and $D$ is the degree matrix of $A$. $\beta$ is the anisotropic diffusion parameter used by the diffusion operator, whose value is empirically set to 0.5 (Vos de Wael et al., 2020). Finally, the functional gradients of multiple components were obtained by calculating the eigenvalues and eigenmatrices of $P_{\beta}$.
The first $N_{gra}$ components of the functional gradients of each ROI were taken as input, and the unsupervised algorithm $K$-means was used to complete the clustering of ROIs. In this way, 116 ROIs were clustered into $k$ communities according to functional similarity. Then we customized a one-dimensional convolution for each community to extract the ROIs features in this community as a representation of this community. The features extraction of each community can be summarized in the following formula.
@@
Table 2: Detailed architecture and parameters of community-clustering scale and whole-communities scale.
Block Layer Output size Explanation
Conv1Di Input (Ni, 64) Ni stands for the number of ROIs belonging to community i
Conv1d (1, 64) outchannel=1,kernel_size=15,padding=’same’,stride=1
LayerNorm (1, 64)
Activation (1, 64)
Dropout (1, 64) dropout rate=0.3
Concatenate (k, 64) The features of k communities are concatenated to form a community network
Conv2D unsqueeze (1, k, 64)
Conv2d (16, 1, 1) outchannel=16,kernelsize=(k, 64),stride=1
squeeze (16)
Linear Linear (16) in features=16,out features=16
Activation (16) At this point, we have obtained a feature representation of the whole brain
2.3.3 The whole-communities scale and classification
By extracting community-clustering scale features, we obtained a community-level network with the shape $(k,64)$. In order to further extract high-level features, Conv2D was used to extract global integration information on whole-communities scale. The parameter settings of Conv2D are listed in Table 2.
In order to train the model serving as feature extractor, after the whole-communities scale module, we used a liner layer and a softmax layer to achieve the supervised classification task. After the model was trained, we used it as feature extractor (denoted as $f(\cdot)$), the outputs of the whole-communities scale module were used as the features.
To implement the classification, the prototype learning was used, which achieves classification by learning a set of prototypes, representing the centers of different classes (Qiu et al., 2024). Prototype learning using extracted multi-perspective brain representations:

### v2

@@ Abstract
Autism Spectrum Disorder (ASD) is a pervasive developmental disorder of the central nervous system, primarily manifesting in childhood. It is characterized by atypical and repetitive behaviors. Currently, diagnostic methods mainly rely on questionnaire surveys and behavioral observations, which are prone to misdiagnosis due to their subjective nature. With advancements in medical imaging, MR imaging-based diagnostics have emerged as a more objective alternative. In this paper, we propose a Hierarchical Neural Network model for ASD identification, termed ASD-HNet, which hierarchically extracts features from functional brain networks based on resting-state functional magnetic resonance imaging (rs-fMRI) data. This hierarchical approach enhances the extraction of brain representations, improving diagnostic accuracy and aiding in the identification of brain regions associated with ASD. Specifically, features are extracted at three levels: (1) the local region of interest (ROI) scale, (2) the community scale, and (3) the global representation scale. At the ROI scale, graph convolution is employed to transfer features between ROIs. At the community scale, functional gradients are introduced, and a K-Means clustering algorithm is applied to group ROIs with similar functional gradients into communities. Features from ROIs within the same community are then extracted to characterize the communities. At the global representation scale, we extract global features from the whole community-scale brain networks to represent the entire brain. We validate the effectiveness of our method using the publicly available Autism Brain Imaging Data Exchange I (ABIDE-I) dataset. Experimental results demonstrate that ASD-HNet outperforms existing methods. The code is available at https://github.com/LYQbyte/ASD-HNet.
@@ 1 Introduction
Autism spectrum disorder (ASD) is a type of pervasive developmental disorder affecting the central nervous system (Pandolfi et al., 2018), approximately 1% of the world’s population is affected by ASD. Moreover, ASD is associated with a high incidence of concurrent disorders, such as depression and anxiety (Vohra et al., 2017). Autism spectrum disorder typically manifests in childhood, with its associated impairments often persisting throughout an individual’s lifetime. Many individuals with ASD require lifelong support to manage their condition (Lord et al., 2018). ASD is characterized by a broad spectrum of symptoms, including deficits in social communication and the presence of restrictive, repetitive behaviors (Hirota and King, 2023; Bhat et al., 2014). The disorder poses significant challenges to both the physical and mental health of affected individuals, profoundly impacting their quality of life as well as that of their families. Despite its profound effects, the precise pathogenesis of ASD remains elusive, making accurate diagnosis particularly challenging.
Unfortunately, there are no definitive drugs and methods for the treatment of ASD, and the pathological mechanisms of ASD remain for further exploring (Geschwind, 2009). At present, the most effective strategies involve providing advanced behavioral training for individuals with ASD, implementing early intervention programs, and addressing problematic behaviors associated with the condition. These approaches underscore the critical importance of early and accurate diagnosis of ASD. However, relying solely on the manifestation of behavioral symptoms often results in delayed diagnosis, potentially causing patients to miss the optimal window for intervention and complicating treatment efforts. Currently, the majority of diagnostic methods rely on questionnaires and behavioral observations to determine whether an individual has ASD (American Psychiatric Association et al., 2013). These approaches are highly subjective and susceptible to environmental influences, which increases the risk of misdiagnosis (Wang et al., 2022). Consequently, there is an urgent need to develop highly accurate, pathology-based diagnostic tools and to identify reliable biomarkers associated with ASD to improve diagnostic precision and facilitate timely intervention.
@@
Benefiting from the rapid development of medical image technology, imaging-based approaches provides us with important tools to analyze and study neurophysiology. Functional magnetic resonance imaging (fMRI), as a non-invasive neuroimaging technique (Matthews and Jezzard, 2004), can measure the blood oxygen level dependent (BOLD) changes caused by neuronal activity throughout the brain at a series of time points (Friston et al., 1994). Numerous studies have demonstrated the feasibility of utilizing resting-state functional magnetic resonance imaging (rs-fMRI) to investigate the interactions between regions of interest (ROIs) in the brains of individuals with mental disorders (Dvornek et al., 2018). In recent years, a growing body of research has integrated deep learning techniques with rs-fMRI to explore and analyze mental illnesses, achieving notable progress (Dong et al., 2020; Luppi et al., 2022). In most studies, the brain is partitioned into multiple ROIs using predefined brain atlases, followed by the extraction of BOLD time series from these ROIs. Subsequently, the correlations between the BOLD time series of different ROIs are computed to construct brain functional connectivity networks (FCNs), which are then used for further analysis and exploration of brain function (Dong and Sun, 2024; Wang et al., 2023; Hong et al., 2019).
@@
Previous deep learning models often lack the model structure exploration of physiological mechanisms and fail to effectively mine multi-scale features, ranging from local regions to the whole-brain scale, which ultimately leads to suboptimal diagnostic performance. In the current study, we combine resting-state functional magnetic resonance imaging with deep learning techniques to design a network model from a brain physiology perspective, with the goal of improving ASD detection performance. We propose a hybrid neural network model for ASD identification, termed ASD-HNet, which hierarchically extracts features from functional brain networks at three scales: local ROI scale, community scale, and global representation scale. Specifically, we use the Automated Anatomical Labeling (AAL) template (Tzourio-Mazoyer et al., 2002) to divide the brain into $N_{R}$ ($N_{R}$=116) regions of interest, extract the BOLD signals from each ROI, and compute the functional connectivity network. A graph convolutional network (GCN) (Niepert et al., 2016) is then employed to transfer information between ROIs. Next, we introduce functional gradients to automatically cluster ROIs into communities, as opposed to relying on pre-defined atlases. This clustering approach groups ROIs with similar functional patterns, allowing us to extract community-scale features. Finally, global brain convolution is applied to the community-scale features to integrate information at the global scale. Additionally, prototype learning is incorporated for classification based on the features extracted by ASD-HNet, leading to improved identification accuracy. Experiments on the publicly available Autism Brain Imaging Data Exchange I (ABIDE-I) dataset validate the superior performance of the proposed ASD-HNet. Moreover, extensive experiments on model generalization further demonstrate the robustness and adaptability of our approach across different datasets.
@@ 2.1 Functional Gradients
Functional gradients in the brain offer a more comprehensive perspective by capturing the continuous spatial patterns of connectivity beyond the segregated networks typically observed in the human brain (Guell et al., 2018; Huntenburg et al., 2018). It’s a gradient-based approach, which performs a nonlinear decomposition of high-dimensional resting-state functional connectivity (rs-FC), can identify distinct brain functional levels by representing connectivity in a continuous, low-dimensional space (Dong et al., 2020). Functional gradients have been extensively applied in brain analysis (Dong et al., 2020; Guo et al., 2023; Gong and Zuo, 2023). In the context of various mental disorders, functional gradients provide a clear indication of abnormal functional hierarchies across brain regions.
@@ 2.2 Materials and data preprocessing
For data preprocessing, we used the data Processing Assistant for Resting-State fMRI (DPARSF) to preprocess rs-fMRI data (Yan and Zang, 2010). The specific steps are as follows: (1) Discarding the first 5 volumes. (2) Correcting of slice-related delay. (3) Correcting head movement. (4) Using EPI template normalization in MNI space, resampling to 3 mm resolution of $3\times 3\times 3$. (5) Smoothing with a 4 mm full-width at half-maximum (FWHM) Gaussian kernel. (6) Linear detrending and temporal band-pass filtering of BOLD signal (0.01-0.10 Hz). (7) Harmful signals that return head movement parameters, white matter, cerebrospinal fluid (CSF), and global signals. The registered fMRI volumes are then partitioned into $N_{R}$ ROIs using the AAL template, and BOLD time series are obtained for each ROI. Based on these time series of all the ROIs, we obtained the brain FCN matrix by calculating the Pearson correlation coefficient with the formula (1).
@@ 2.3 Proposed Method
The schematic diagram of the proposed ASD-HNet is shown in Figure 2. This model is designed to progressively extract the discriminative and robust representations based on FCNs from three different scales, i.e., local ROI scale, community scale, and global representation scale, to achieve ASD identification. More details are described below.
@@ 2.3.1 The local ROI scale
Using the F-score, we derived two sparse FCN matrices. Following the methods in (Zhang et al., 2023a; Tong et al., 2023), we retained the top 10% of connections in the functional connectivity matrix to construct a sparse FCN matrix, which served as the adjacency matrix ($E$) for building a graph network. Additionally, to mitigate noise in the node features, we further sparsified the FCN matrix by retaining a proportion ${\lambda_{R}}$ (${\lambda_{R}}$=50%) based on the F-score criterion.
Based on these two sparse FCN matrices, we constructed a ROI-scale graph network $G(V,E)$, where $V$ is the node feature matrix and $E$ is the adjacency matrix. Then, we used graph convolutional neural network to conduct feature extraction on the ROI scale. Here, we used one layer of spatial-based GCN to complete the message transfer between ROIs, the specific implementation is as follows:
@@
where $W$ is a layer-specific trainable weight matrix with the shape ($N_{R}$, $D_{1}$) ($D_{1}$=64) and $\sigma$ is the activation function. In order to make nodes pay attention to their own characteristics, the diagonal elements of the adjacency matrix $E$ are all set to 1.0.
Following the graph convolution operation, we further employed a ROIs attention mechanism to weight each ROI features. Sepcifically, an attention weight matrix $H_{atten}$ was initialized with all elements set to 1.0. Subsequently, two successive fully connected layers were employed to learn the attention weights for each ROI. The operation of a single linear layer is shown in the following formula (4). It should be noted that since ROIs are weighted using positive numbers, so the activation function ($\sigma_{atten}$) was ReLU in the fully connected layer.
@@
After obtaining the attention weights of ROIs, we applied these weights to multiply the features of each ROI. Subsequently, a residual connection was employed to integrate the weighted ROI features with the original ROI features, followed by a batch normalization operation, as illustrated in formula (5).
@@
2.3.2 The community scale
The brain’s community network comprises groups of functionally similar ROIs, each with a high degree of functional concentration. Community networks are crucial for understanding the functional organization of the brain (Van Den Heuvel et al., 2009). Functional communities have been shown to be relevant to understanding cognitive behavior (Van Den Heuvel and Pol, 2010), mental states (Geerligs et al., 2015), and neurological and psychiatric disorders (Canario et al., 2021). We employed the community scale representation module to further extracted functional features at community-clustering scale as illustrated in Figure 2. Since functional gradients can well reflect the functional similarity between ROIs, we proposed to use functional gradients to automatically cluster ROIs into community networks.
We utilized whole group-level (include ASD subjects and HC subjects) average functional connectivity matrix inputs into BrainSpace (Vos de Wael et al., 2020) to obtain group-level functional gradients. In general, the computation of functional gradients involves two steps. The first step is selecting a kernel function to compute the affinity matrix. In the second step, the resulting affinity matrix is subjected to nonlinear dimensionality reduction to obtain the functional gradients. In our work, we chose cosine similarity as the kernel function to compute the affinity matrix ($A(i,j)$). The calculation is expressed as follows:
@@ 2.3.2 The community scale
where $A$ is the affinity matrix, and $D$ is the degree matrix of $A$. $\beta$ is the anisotropic diffusion parameter used by the diffusion operator, whose value is empirically set to 0.5 (Vos de Wael et al., 2020). Then, the eigenvalue decomposition algorithm is used to extract the top $n$ largest eigenvalues ($\lambda$) along with their corresponding eigenvectors ($v$).
${{P_{\beta}}v=\lambda v}$ (9)
Finally, the diffusion embedding is constructed by scaling the eigenvectors to derive the $i$-th functional gradient component ${\Phi_{i}}$.
${\Phi_{i}}=\frac{{{\lambda_{i}}}}{{1-{\lambda_{i}}}}\cdot{v_{i}}$ (10)
The first $N_{gra}$ ($N_{gra}$=2) components of the functional gradients of each ROI were taken as input, and the unsupervised algorithm $K$-means was used to complete the clustering of ROIs. In this way, $N_{R}$ ROIs were clustered into $k$ ($k$=7) communities according to functional similarity. Then we customized a 1D convolution for each community to extract the ROIs features in this community as a representation of this community. The features extraction of each community can be summarized in the following formula.
@@
Table 2: Detailed architecture and parameters of community scale and global representation scale.
Module Block Layer Output size Explanation
Community Scale Conv1Di Input (Ni, D1) Ni stands for the number of ROIs belonging to community i, D1=64
Conv1d (1, D1) outchannel=1, kernel_size=15, padding=’same’,stride=1
LayerNorm (1, D1)
Activation (1, D1)
Dropout (1, D1) dropout rate=0.3
Concatenate (k, D1) The features of k communities are concatenated to form a community network
Global Scale Conv2D unsqueeze (1, k, D1)
Conv2d (D2, 1, 1) outchannel=16, kernelsize=(k, D1), stride=1, D2=16
squeeze (D2)
Linear Linear (D2) in features=D2, out features=D2
Activation (D2) At this point, we have obtained a feature representation of the whole brain
2.3.3 The global representation scale and classification
By clustering functional gradients and extracting representations of each community, we obtained a functional community network with the shape $(k,D_{1})$. To effectively characterize the brain at a whole-brain scale, we proposed a global scale representation module, which consists of a 2D convolutional layer (with a kernel size equal to the scale of the community network) and a linear layer to achieve whole-brain representation. The parameter details of this module are listed in Table 2.
In order to train the model serving as feature extractor, after the global scale representation module, we used a liner layer and a softmax layer to achieve the supervised classification task (i.e., training process). After the model was trained, we used it as feature extractor (denoted as $f(\cdot)$), the outputs of the global scale representation module were used as the subject’s characteristic features.
To implement the classification, the prototype learning was introduced, which achieves classification by learning a set of prototypes, representing the centers of different classes (Qiu et al., 2024). The prototype of the $i$-th class is computed as follows: